solve for x
4(x+4)=6(y-2)

Answer by checkley71(8403)

4(X+4)=6(Y-2)
4X+16=6Y-12
4X=6Y-12-16
4X=6Y-28
X=6Y/4-28/4
X=1.5Y-7 ANSWER.